OpenStudy (anonymous):

The point (3, -4) is on the terminal edge of an angle theta. Find the exact value of the 6 trigonometric functions for theta. This point creates a 3-4-5 triangle, so wouldn't the arc described by theta be approximate? How do I proceed!?

OpenStudy (anonymous):

Is it just sin = 4/5, cos = 3/5, tan = 4/3 etc?